Problem teos2 SNF 2007-09-13 19:59:33: unstable temperature

panjun at snf.stanford.edu panjun at snf.stanford.edu
Thu Sep 13 19:59:34 PDT 2007


in the deposition this morning, the templ was 250, which was later corrected by staff
in the evening deposition, the temps is 250 at the beginning of step 70, and later jumped to 702.0B.  TCU% is 0 sometimes




More information about the teos2-pcs mailing list